### v2

@@ 1 Introduction
Speaker diarization consists in identifying “who spoke when” in an input audio, by segmenting it into speaker-attributed regions (Anguera et al., 2012; Park et al., 2022) that correspond to speakers’ utterances. It is an essential pre-processing task in many applications, such as lectures, meetings, live captioning, speaker-based indexing, telephone calls, and doctor-patient conversations.
@@
The majority of the methods above only work offline and thus are not suitable for streaming processing. The extension from offline to online processing is way simpler for EEND. Clustering-based systems consist of a pipeline of several modules (i.e., voice activity detection, speaker embedding extraction, clustering, etc). Each of these modules has to be updated in order to work online. In contrast, end-to-end methods can be easily adapted for online diarization by employing a speaker-tracing buffer (STB) to store previous input-output pairs that can be exploited for online inference (Xue et al., 2021a, b; Horiguchi et al., 2023). An alternative strategy is to simply replace the neural architecture with another one that allows low-latency streaming processing (Han et al., 2021). A similar approach, which is the basis for this work, is proposed by Morrone et al. (2022a) also for SSGD.
In this preliminary work, SSep and VAD models were trained independently and combined together with no additional training. Despite its simplicity, this approach only reaches a sub-optimal diarization performance. Additionally, a major disadvantage is that the SSep module needs a dataset in which oracle sources for the speakers are available. This could lead to mismatched training-inference conditions as in most real-world scenarios oracle speaker sources are difficult to obtain (Subakan et al., 2022). In this paper, we build upon this previous work (Morrone et al., 2022a) and propose a fully end-to-end integration of the SSep and VAD modules. This approach only requires diarization labels during the fine-tuning stage and thus eliminates the need for oracle speaker target sources. We show that the proposed end-to-end strategy provides significant improvements on two widely used datasets, i.e., Fisher Corpus (Cieri et al., 2004) and CALLHOME (Przybocki and Alvin, 2001). These reflect two different training vs. inference conditions: fully matched and unmatched respectively. In particular, our best online model outperforms the current state-of-the-art EEND system on the $2$-speaker subset of the CALLHOME dataset despite having an order of magnitude lower latency, i.e., $0.1$ vs $1$ seconds. Moreover, we review and extend our previous work (Morrone et al., 2022a) with additional analysis. The other contributions with respect to the previous paper are summarized below:
@@
  • We thoroughly study the effect of the leakage removal algorithm on both disjoint and end-to-end trained SSGD.
@@ 1.1 Related Works
Recently, deep learning methods are employed for end-to-end neural diarization (EEND) approaches. A major advantage of EEND is that they are able to deal with overlapped speech without any modification. The first EEND methods perform diarization as a simple multi-speaker voice activity detection (VAD) problem, in which each output represents a different speaker’s speech activity. EEND systems can employ different neural architectures, such as bidirectional long-short memory (Fujita et al., 2019a) and self-attention (SA-EEND) (Fujita et al., 2019b). EEND-based systems are trained directly to perform diarization using permutation invariant training (PIT) (Kolbæk et al., 2017) as the diarization problem is inherently permutation-invariant without any a-priori information. With enough training data, end-to-end approaches have been shown to outperform current state-of-the-art clustering-based systems (Horiguchi et al., 2021c). Contrary to clustering-based approaches, the maximum number of total speakers is fixed in the aforementioned EEND architectures. Additionally, end-to-end systems need to process the entire input signal during inference, resulting in significant memory consumption for long recordings (e.g., >10 minutes). Chunk-wise processing can help but is not viable as it leads to speaker label permutation across chunks due to PIT. Horiguchi et al. (2020) solves the first problem by extending the basic EEND with an auto-regressive encoder-decoder (EEND-EDA) architecture. The second issue is generally addressed by combining clustering and EEND. In Horiguchi et al. (2021a) EEND is exploited to refine the results of a clustering-based algorithm. Bredin and Laurent (2021) proposes a speaker segmentation model inspired by EEND to perform overlap-aware resegmentation in a conventional diarization pipeline. However, these approaches result in more complex systems in contrast to the simplicity of fully end-to-end architectures. A tighter integration of EEND and clustering, i.e., EEND-vector clustering (EEND-VC), is proposed in Kinoshita et al. (2021a, b). It performs chunk-wise processing to ensure that a given maximum number of active speakers can be present in each chunk (e.g., 2 or 3). The original EEND is modified to output global speaker embeddings that are aggregated across chunks using a constrained clustering algorithm. This method can both deal with an arbitrary number of speakers and solve the inter-chunk speaker permutation problem. An approach similar to EEND-VC, named EEND with global and local attractors (EEND-GLA), is proposed in Horiguchi et al. (2021b) which combines EEND-EDA and unsupervised clustering to deal with cases where the number of speakers appearing during inference is higher than that during training. Another system (Zeghidour et al., 2021) employs a different architecture that iteratively builds embeddings for each speaker which are exploited to condition a VAD module.
@@
An alternative framework to deal with overlapped speech is continuous speech separation (CSS) (Chen et al., 2020b; Morrone et al., 2022b). CSS extends PIT-based SSep to long recording scenarios, by applying separation in a chunk-wise manner, where each chunk is assumed to contain a fixed number of speakers (usually 2-3). Since the underlying separator is trained via a PIT objective, output permutation consistency between chunks is not guaranteed. CSS solves this problem by performing overlapping inference (i.e., using strides shorter than chunk sizes) and reordering adjacent chunks based on a similarity measure over the portion in which they overlap. Several recent works have proposed diarization systems inspired by CSS. In  Raj et al. (2021a) and Xiao et al. (2021) separation is done in windowed segments. In this case, speaker permutation is addressed by applying a diarization method (e.g., clustering-based) across the separated audio streams. On the other hand, Fang et al. (2021) proposes speech separation guided diarization (SSGD), where diarization is performed by first separating the input mixture and applying a conventional VAD to detect speech segments in each channel. In Morrone et al. (2022a), we improve the SSGD architecture using a neural-based VAD and a novel post-processing algorithm that removes the channel leakage generated by separation. In addition, SSGD is adapted to allow online inference employing causal SSep and VAD models.
@@
All the aforementioned approaches, with the exception of Morrone et al. (2022a), are not suitable for streaming applications (e.g., live captioning) as they only work offline. To deal with online processing, Xue et al. (2021b) extends the SA-EEND with an STB mechanism. Inference is performed on short chunks and the speaker permutation information is selected from input-output pairs of previous frames and stored in a buffer. The permutation ambiguity is solved using the buffered frames to condition the output of the current chunk. Xue et al. (2021a) and Horiguchi et al. (2023) propose similar STB-based extensions for EEND-EDA and EEND-GLA, respectively. Instead, Han et al. (2021) improves the online EEND-EDA using chunk-level recurrence to process the chunk hidden states making the model complexity linear in time. Coria et al. (2021) designs a different approach that combines the use of EEND with an x-vector extractor and online clustering. The EEND model is used to gate the representation before the x-vector statistical pooling layer, to extract per-speaker embeddings even in overlap regions.
@@ 2.1 Speech Separation Module
In our experiments we employ causal separation models as SSep modules (i.e., Conv-TasNet (Luo and Mesgarani, 2019), DPTNet (Chen et al., 2020a) and DPRNN (Luo et al., 2020)). A streaming modified version of DPTNet is attained in this work by simply masking the future frames in each self-attention layer in the inter-processing blocks, with the rest being equal to causal DPRNN (DPTNet is identical to DPRNN except for self-attention layers).
To compare the proposed SSGD with both clustering-based and EEND state-of-the-art offline systems, we also experiment with non-causal SSep models. Additionally, we consider the use of CSS with non-causal SSep models. In such a configuration, the latency of these models is tied to the CSS window size and thus making offline SSep models suitable for streaming inference. We do not apply CSS in causal SSep models as they do not require chunk-wise processing to work online.
@@ 2.3 Leakage Removal Module
The leakage removal post-processing algorithm is summarized in Algorithm 1. The algorithm is fed with an input mixture $\mathbf{Y}$ and two estimated sources $\mathbf{X}^{1}$ which are split into disjoint segments $\mathbf{Y}_{\ell}$, $\mathbf{X}_{\ell}^{1}$, $\mathbf{X}_{\ell}^{2}$ of length $L$. For each segment, we compute the Scale-Invariant Signal-to-Distortion Ratios (SI-SDR) (Le Roux et al., 2019) $s_{\ell}^{1}$, $s_{\ell}^{2}$ between segments of every source $\mathbf{X}_{\ell}^{1}$, $\mathbf{X}_{\ell}^{2}$ with the associated segment $\mathbf{Y}_{\ell}$ of input mixture. A leaked segment is detected when both $s_{\ell}^{1}$, $s_{\ell}^{2}$ are above a threshold $t_{\ell r}$. In the SSGD framework leakage is removed by filling with zeros the segments with lower SI-SDR. Then, the post-processed estimated sources $\mathbf{\tilde{X}_{\ell}}$ are passed as input to the following VAD module.
@@ 2.4 End-to-End Training
In the end-to-end SSGD SSep and VAD are jointly optimized, while removing the leakage removal during training. The SSep and VAD modules are initialized with the parameters of the disjoint SSGD and fine-tuned following two methods. In the first approach, i.e., VAD fine-tuning, we freeze the SSep model and only optimize the VAD parameters. Instead, in the SSep+VAD fine-tuning method all parameters are optimized jointly.
Also for end-to-end fine-tuned SSGD systems we apply the proposed leakage removal post-processing algorithm. However, for end-to-end SSGD we apply the masking directly on the VAD output to avoid changing the distribution of estimated sources. This is because we found that doing so could degrade the VAD performance after end-to-end adaptation, since it was fine-tuned on the output of the SSep module.
@@ 3.2 Architecture, Training and Inference Details
We experiment with SSep architectures both in online and offline settings. In particular, we consider $3$ different SSep models: Conv-TasNet, DPTNet and DPRNN. For Conv-TasNet we employ the best hyperparameter configuration proposed by Luo and Mesgarani (2019). Instead, for DPRNN and DPTNet we use the hyperparameters proposed in Luo et al. (2020) and Chen et al. (2020a), respectively, with some changes. The kernel size for encoder/decoder is increased to $16$ to reduce memory consumption. The chunk and hop sizes are set to $100$ and $50$ (i.e., $50$% overlap). In addition, the global layer normalization is replaced with standard layer normalization for online/causal models. We use the implementations available through the Asteroid toolkit (Pariente et al., 2020). We pre-train SSep models on the simulated fully overlapped Fisher maximizing the SI-SDR function. We observed that the SSep pre-training converged much faster compared to training from scratch on real data. We use Adam optimizer (Kingma and Ba, 2015), batch size $4$ and learning rate $10^{-3}$. We clip gradients with $l_{2}$ norm greater than $5$. Then, we fine-tune each SSep model on $60$ s long random segments from real Fisher recording using a batch size of $1$ and reducing the learning rate to $10^{-4}$. For Conv-TasNet and DPRNN, learning rate is halved whether SI-SDR does not improve on the validation set for $10$ consecutive epochs. For DPTNet, we employ the learning rate scheduler used in Chen et al. (2020a) with $k_{1}=1$, $k_{2}=4\cdot 10^{-4}$, $d_{model}=64$, and $warmup\_n=5n_{epoch}$, where $n_{epoch}$ is the number of training steps performed in a single epoch. If no improvement is observed for $20$ epochs on validation, training is stopped. For the CALLHOME dataset, the separators are adapted on the simulated CALLHOME using the same hyperparameters of the fine-tuned models, except for the length of training segments which is set to $30$ s.
@@
Method VAD Leakage Removal Latency (s) MS FA SC DER SA-EEND w/STB (Xue et al., 2021b) n.a. n.a. 1 12.5 BW-EDA-EEND (Han et al., 2021) 10 11.8 SA-EEND-EDA w/STB (Xue et al., 2021a) 10 10.0 EEND-GLA w/BW-STB (Horiguchi et al., 2023) 1 9.0 Conv-TasNet Energy 0.01 7.4 35.3 3.0 45.7 DPTNet 0.1 5.5 6.3 0.8 12.6 DPRNN 0.1 5.7 5.2 1.8 12.7 Conv-TasNet TCN 0.01 12.2 36.2 5.1 53.4 DPTNet 0.1 6.6 4.5 0.6 11.7 DPRNN 0.1 5.9 3.8 1.7 11.6 Conv-TasNet TCN 0.01 13.2 6.3 13.0 32.5 DPTNet 0.1 6.3 2.6 1.2 10.0 DPRNN 0.1 6.8 2.2 2.2 11.2
Table 2: Disjoint SSGD: diarization results on the CALLHOME test set in the online scenario. Diarization is assessed using diarization error rate (DER), missed speech (MS), false alarm (FA) and speaker confusion errors (SC). Latency of the system is reported in seconds. The best results among proposed techniques are shown in bold, and among EEND methods are underlined.
The VAD is trained on $2$ s long segments and the batch size is set to $256$. We employ the same optimizer, learning rate scheduler, gradient clipping and early stopping policy used for Conv-TasNet and DPRNN. At inference the VAD is applied on estimated separated sources independently. For each frame, the VAD predictions above a threshold $t_{v}$ are labeled as speech. Then, the thresholded predictions are smoothed using a median filter and segments shorter than a threshold $t_{s}$ are removed to mitigate false alarm errors. The median filter length, $t_{v}$ and $t_{s}$ parameters are tuned on the Fisher validation set for each SSGD architecture. Likewise, these parameters are tuned on the CALLHOME adaptation set for CALLHOME models. The threshold $t_{\ell r}$ and the segment length $L$ of the leakage removal algorithm are tuned manually and set to $3$ and $0.1$ s, respectively, to not affect the latency of the VAD.

### v3

@@ 1 Introduction
Speaker diarization consists in identifying “who spoke when” in an input audio, by segmenting it into speaker-attributed regions (Anguera et al., 2012; Park et al., 2022; Serafini et al., 2023) that correspond to speakers’ utterances. It is an essential pre-processing task in many applications, such as lectures, meetings, live captioning, speaker-based indexing, telephone calls, and doctor-patient conversations.
@@
The majority of the methods above only work offline and thus are not suitable for streaming processing. The extension from offline to online processing is way simpler for EEND. Clustering-based systems consist of a pipeline of several modules (i.e., VAD, speaker embedding extraction, clustering, etc). Each of these modules has to be updated in order to work online. In contrast, end-to-end methods can be easily adapted for online diarization by employing a speaker-tracing buffer (STB) to store previous input-output pairs that can be exploited for online inference (Xue et al., 2021a, b; Horiguchi et al., 2023). An alternative strategy is to simply replace the neural architecture with another one that allows low-latency streaming processing (Han et al., 2021). A similar approach, which is the basis for this work, is proposed by Morrone et al. (2022a) also for SSGD.
In this preliminary work, SSep and VAD models were trained independently and combined together with no additional training. Despite its simplicity, this approach only reaches a sub-optimal diarization performance. Additionally, a major disadvantage is that the SSep module needs a dataset in which oracle sources for the speakers are available. This could lead to mismatched training-inference conditions as in most real-world scenarios oracle speaker sources are difficult to obtain (Subakan et al., 2022). In this paper, we build upon our previous work (Morrone et al., 2022a) and propose a fully end-to-end integration of the SSep and VAD modules. This approach only requires diarization labels during the fine-tuning stage and thus eliminates the need for oracle speaker target sources. We focus on the conversational telephone speech (CTS) domain, where the maximum number of speakers is limited to $2$. Despite this limitation, this scenario is quite common in many commercial applications (e.g., doctor-patient recordings). This choice also allows to compare directly with previous works on EEND diarization. We show that the proposed end-to-end strategy provides significant improvements on two widely used datasets, i.e., Fisher Corpus (Cieri et al., 2004) and CALLHOME (Przybocki and Alvin, 2001). These reflect two different training vs. inference conditions: fully matched and unmatched respectively. In particular, our best online model outperforms the current state-of-the-art EEND system on the $2$-speaker subset of the CALLHOME dataset despite having an order of magnitude lower latency, i.e., $0.1$ vs $1$ seconds.
The main contributions with respect to our previous work (Morrone et al., 2022a) are summarized below:
  • We propose a joint fine-tuning strategy of SSep and VAD modules which consistently improves over disjoint trained SSGD. Notably, the proposed approach allows fine-tuning on datasets for which separated oracle sources are not available.
  • We thoroughly study the effect of the proposed leakage removal algorithm on both disjoint and end-to-end trained SSGD. Additionally, we analyze the relationship between the leakage removal aggressiveness and the diarization evaluation metrics.
@@
Although it is straightforward to cascade causal SSep and VAD modules to achieve low latency, our experiments show that simple concatenation does not lead to optimal performance. Firstly, the choice of appropriate SSep and VAD models is crucial. Moreover, the use of the leakage removal algorithm consistently improves accuracy for all proposed methods at a very negligible cost. Finally, joint fine-tuning further reduces diarization errors and it is particularly appealing when oracle separated sources are not available (e.g., CALLHOME). We complete our work providing interesting in-depth analysis on several SSGD aspects (e.g., online/offline systems comparison, latency analysis, leakage removal impact, ASR evaluation). Such analysis aims at shedding more light on the pros and cons of using speech separation for the diarization task.
@@ 1.1 Related Works
Recently, deep learning methods are employed for end-to-end neural diarization approaches. A major advantage of EEND is that they are able to deal with overlapped speech without any modification. The first EEND methods perform diarization as a simple multi-speaker voice activity detection problem, in which each output represents a different speaker’s speech activity. EEND systems can employ different neural architectures, such as bidirectional long-short memory (Fujita et al., 2019a) and self-attention (SA-EEND) (Fujita et al., 2019b). EEND-based systems are trained directly to perform diarization using permutation invariant training (PIT) (Kolbæk et al., 2017) as the diarization problem is inherently permutation-invariant without any a-priori information. With enough training data, end-to-end approaches have been shown to outperform current state-of-the-art clustering-based systems (Horiguchi et al., 2021c). Contrary to clustering-based approaches, the maximum number of total speakers is fixed in the aforementioned EEND architectures. Additionally, end-to-end systems need to process the entire input signal during inference, resulting in significant memory consumption for long recordings (e.g., >10 minutes). Chunk-wise processing can help but is not viable as it leads to speaker label permutation across chunks due to PIT. Horiguchi et al. (2020) solves the first problem by extending the basic EEND with an auto-regressive encoder-decoder (EEND-EDA) architecture. The second issue is generally addressed by combining clustering and EEND. In Horiguchi et al. (2021a) EEND is exploited to refine the results of a clustering-based algorithm. Bredin and Laurent (2021) proposes a speaker segmentation model inspired by EEND to perform overlap-aware resegmentation in a conventional diarization pipeline. However, these approaches result in more complex systems in contrast to the simplicity of fully end-to-end architectures. A tighter integration of EEND and clustering, i.e., EEND-vector clustering (EEND-VC), is proposed in Kinoshita et al. (2021a, b). It performs chunk-wise processing to ensure that a given maximum number of active speakers can be present in each chunk (e.g., 2 or 3). The original EEND is modified to output global speaker embeddings that are aggregated across chunks using a constrained clustering algorithm. This method can both deal with an arbitrary number of speakers and solve the inter-chunk speaker permutation problem. An approach similar to EEND-VC, named EEND with global and local attractors (EEND-GLA), is proposed in Horiguchi et al. (2021b) which combines EEND-EDA and unsupervised clustering to deal with cases where the number of speakers appearing during inference is higher than that during training. Another system (Zeghidour et al., 2021) employs a different architecture that iteratively builds embeddings for each speaker which are exploited to condition a VAD module.
@@
An alternative framework to deal with overlapped speech is continuous speech separation (CSS) (Chen et al., 2020b; Morrone et al., 2022b). CSS extends PIT-based SSep to long recording scenarios, by applying separation in a chunk-wise manner, where each chunk is assumed to contain a fixed number of speakers (usually 2-3). Since the underlying separator is trained via a PIT objective, output permutation consistency between chunks is not guaranteed. CSS solves this problem by performing overlapping inference (i.e., using strides shorter than chunk sizes) and reordering adjacent chunks based on a similarity measure over the portion in which they overlap. Several recent works have proposed diarization systems inspired by CSS. In  Raj et al. (2021a) and Xiao et al. (2021) separation is done in windowed segments. In this case, speaker permutation is addressed by applying a diarization method (e.g., clustering-based) across the separated audio streams. On the other hand, Fang et al. (2021) proposes SSGD, where diarization is performed by first separating the input mixture and applying a conventional VAD to detect speech segments in each channel. In Morrone et al. (2022a), we improve the SSGD architecture using a neural-based VAD and a novel post-processing algorithm that removes the channel leakage generated by separation. In addition, SSGD is adapted to allow online inference employing causal SSep and VAD models.
@@
All the aforementioned approaches are not suitable for streaming applications (e.g., live captioning) as they only work offline. Although one could potentially leverage CSS-based systems to allow online inference (as done in Yoshioka et al. (2019)), no established solutions that exploit low-latency SSep for diarization are available in literature, with the exception of Morrone et al. (2022a). Specifically, online inference is not trivial for systems that integrate CSS and clustering-based diarization (Raj et al., 2021a; Xiao et al., 2021) as all submodules have to work online. Several methods have been proposed to allow EEND systems to deal with online processing. Xue et al. (2021b) extends the SA-EEND with an STB mechanism. Inference is performed on short chunks and the speaker permutation information is selected from input-output pairs of previous frames and stored in a buffer. The permutation ambiguity is solved using the buffered frames to condition the output of the current chunk. Xue et al. (2021a) and Horiguchi et al. (2023) propose similar STB-based extensions for EEND-EDA and EEND-GLA, respectively. Instead, Han et al. (2021) improves the online EEND-EDA using chunk-level recurrence to process the chunk hidden states making the model complexity linear in time. Coria et al. (2021) designs a different approach that combines the use of EEND with an x-vector extractor and online clustering. The EEND model is used to gate the representation before the x-vector statistical pooling layer, to extract per-speaker embeddings even in overlap regions.
@@ 2.1 Speech Separation Module
In our experiments we employ causal separation models as SSep modules (i.e., Conv-TasNet (Luo and Mesgarani, 2019), DPTNet (Chen et al., 2020a) and DPRNN (Luo et al., 2020)). To compare the proposed SSGD with both clustering-based and EEND state-of-the-art offline systems, we also experiment with non-causal SSep models.
Low latency is achieved in different ways according to the SSep architectures. For Conv-TasNet, we only use causal convolutions and global layer normalization is replaced with cumulative layer normalization. For DPRNN, online processing is achieved using LSTM in place of bi-directional LSTM for inter-chunk processing to tie the latency to the intra-chunk segment size (Li and Luo, 2023). Regarding DPTNet, a streaming modified version is attained in this work by simply masking the future frames in each self-attention layer in the inter-processing blocks, with the rest being equal to causal DPRNN (DPTNet is identical to DPRNN except for self-attention layers).
Additionally, we consider an alternative approach that uses non-causal SSep models in the CSS framework to allow streaming inference (as done in Chen et al. (2020b) and Morrone et al. (2022b)). In such a configuration, the latency of non-causal models is tied to the CSS window size. We do not apply CSS in causal SSep models as they do not require chunk-wise processing to work online.
@@ 2.3 Leakage Removal Module
The leakage removal post-processing algorithm is summarized in Algorithm 1. The algorithm is fed with an input mixture $\mathbf{Y}$ and two estimated sources $\mathbf{X}^{1}$, $\mathbf{X}^{2}$ which are split into disjoint segments $\mathbf{Y}_{\ell}$, $\mathbf{X}_{\ell}^{1}$, $\mathbf{X}_{\ell}^{2}$ of length $L$. For each segment, we compute the Scale-Invariant Signal-to-Distortion Ratios (SI-SDR) (Le Roux et al., 2019) $s_{\ell}^{1}$, $s_{\ell}^{2}$ between segments of every source $\mathbf{X}_{\ell}^{1}$, $\mathbf{X}_{\ell}^{2}$ with the associated segment $\mathbf{Y}_{\ell}$ of input mixture. A leaked segment is detected when both $s_{\ell}^{1}$, $s_{\ell}^{2}$ are above a threshold $t_{\ell r}$. In the SSGD framework leakage is removed by filling with zeros the segments with lower SI-SDR. Then, the post-processed estimated sources $\mathbf{\tilde{X}_{\ell}}$ are passed as input to the following VAD module.
@@ 2.4 End-to-End Training
In the end-to-end SSGD SSep and VAD are jointly optimized, while disabling the leakage removal during training. The SSep and VAD modules are initialized with the parameters of the disjoint SSGD and fine-tuned following two methods. In the first approach, i.e., VAD fine-tuning, we freeze the SSep model and only optimize the VAD parameters. Instead, in the SSep+VAD fine-tuning method all parameters are optimized jointly.
For end-to-end trained SSGD systems we apply a modified version of the proposed leakage removal post-processing algorithm. We found that the VAD performance could degrade when Algorithm 1 is applied as it is. Indeed, the VAD module is fine-tuned on the output of the SSep module, then changing the separated sources during inference introduces a mismatch with training conditions. In this case, we only employ the Algorithm 1 to detect leaked segments in separated estimated sources without filling them with zeros. At the end, VAD output frames associated to leaked segments are explicitly marked as non-speech frames.
@@ 3.2 Architecture, Training and Inference Details
We experiment with SSep architectures both in online and offline settings. In particular, we consider $3$ different SSep models: Conv-TasNet, DPTNet and DPRNN. For Conv-TasNet we employ the best hyperparameter configuration proposed by Luo and Mesgarani (2019). Instead, for DPRNN and DPTNet we use the hyperparameters proposed in Luo et al. (2020) and Chen et al. (2020a), respectively, with some changes. The kernel size for encoder/decoder is increased to $16$ to reduce memory consumption. The chunk and hop sizes are set to $100$ and $50$ (i.e., $50$% overlap). In addition, the global layer normalization is replaced with standard layer normalization for online/causal models. We use the implementations available through the Asteroid toolkit (Pariente et al., 2020). We pre-train SSep models on the simulated fully overlapped Fisher maximizing the SI-SDR function. We observed that the SSep pre-training allowed faster convergence compared to training from scratch on real Fisher data. We use Adam optimizer (Kingma and Ba, 2015), batch size $4$ and learning rate $10^{-3}$. We clip gradients with $l_{2}$ norm greater than $5$. Then, we continue the training of each SSep model on $60$ s long random segments from real Fisher recording using a batch size of $1$ and reducing the learning rate to $10^{-4}$. For Conv-TasNet and DPRNN, learning rate is halved whether SI-SDR does not improve on the validation set for $10$ consecutive epochs. For DPTNet, we employ the learning rate scheduler used in Chen et al. (2020a) with $k_{1}=1$, $k_{2}=4\cdot 10^{-4}$, $d_{model}=64$, and $warmup\_n=5n_{epoch}$, where $n_{epoch}$ is the number of training steps performed in a single epoch. If no improvement is observed for $20$ epochs on validation, training is stopped. For the CALLHOME dataset, the separators are adapted on the simulated CALLHOME using the same hyperparameters of the fine-tuned models, except for the length of training segments which is set to $30$ s.
@@
The VAD is trained on $2$ s long segments and the batch size is set to $256$. We employ the same optimizer, learning rate scheduler, gradient clipping and early stopping policy used for Conv-TasNet and DPRNN. At inference the VAD is applied on estimated separated sources independently. For each frame, the VAD predictions above a threshold $t_{v}$ are labeled as speech. Then, the thresholded predictions are smoothed using a median filter and segments shorter than a threshold $t_{s}$ are removed to mitigate false alarm errors. The median filter length, $t_{v}$ and $t_{s}$ parameters are tuned on the Fisher validation set for each SSGD architecture. Likewise, these parameters are tuned on the CALLHOME adaptation set for CALLHOME models. The threshold $t_{\ell r}$ and the segment length $L$ of the leakage removal algorithm are manually tuned and set to $3$ and $0.01$ s, respectively, so as not to affect the latency of the VAD.
For SSGD methods, algorithmic latencies of all modules correspond to their respective online processing unit lengths. The online processing unit is the minimum amount of buffered data needed to produce new outputs. It represents the ideal, lower bound latency that can be obtained if the computational processing times of all modules are zero. In this paper, we adopt this convention to compare latency across different diarization systems (Xue et al., 2021b; Han et al., 2021; Xue et al., 2021a; Horiguchi et al., 2023). For speech enhancement and separation algorithms the online processing unit often corresponds to the STFT (Wang and Watanabe, 2022) or learned filterbank synthesis windows (Luo and Mesgarani, 2018, 2019). On the other hand, for separators based on the dual-path architecture (e.g., DPRNN and DPTNet) the online processing unit length is decided by the intra-chunk segment size (Luo et al., 2020; Li et al., 2022). Since latency is tied to the online processing unit length, the total algorithmic latency is equal to the maximum latency among all latencies of the modules of the SSGD pipeline. However, being the modules connected in series, real-world latency on actual hardware would instead be the sum between the total algorithmic latency and the processing times sum of all modules. Actual latency times that can been obtained under specific hardware conditions are reported in Section 4.7. For systems based on DPRNN and DPTNet the latency is set to $0.1$ s, which is the latency of the separators. In contrast, Conv-TasNet has lower latency than the VAD and the leakage removal algorithm, as such the resulting latency is dictated by the VAD at $0.01$ s. Full details about latency computation can be found in Appendix A.
@@
Method VAD Leakage Removal Latency (s) MS FA SC DER SA-EEND w/STB (Xue et al., 2021b) n.a. n.a. 1 12.5 BW-EDA-EEND* (Han et al., 2021) 10 11.8 SA-EEND-EDA w/STB* (Xue et al., 2021a) 10 10.0 EEND-GLA w/BW-STB* (Horiguchi et al., 2023) 1 9.0 Conv-TasNet Energy 0.01 7.4 35.3 3.0 45.7 DPTNet 0.1 5.5 6.3 0.8 12.6 DPRNN 0.1 5.7 5.2 1.8 12.7 Conv-TasNet TCN 0.01 12.2 36.2 5.1 53.4 DPTNet 0.1 6.6 4.5 0.6 11.7 DPRNN 0.1 5.9 3.8 1.7 11.6 Conv-TasNet TCN 0.01 13.2 6.3 13.0 32.5 DPTNet 0.1 6.3 2.6 1.2 10.0 DPRNN 0.1 6.8 2.2 2.2 11.2
Table 2: Disjoint SSGD: diarization results on the CALLHOME test set in the online scenario. Diarization is assessed using diarization error rate (DER), missed speech (MS), false alarm (FA) and speaker confusion errors (SC). Algorithmic latency is reported in seconds. The best results among proposed techniques are shown in bold, and among EEND methods are underlined.
*The number of speakers in input recordings is estimated by the model.